OpenStudy (anonymous):

Would 3r + 8 = -16 be -8?

OpenStudy (lgbasallote):

3r + 8 = -18 3r = -16 - 8 3r = -24 r = -8 wow! you're right! congrats :D

OpenStudy (shubhamsrg):

yep..the very same..

OpenStudy (lgbasallote):

the first line is 3r + 8 = -16 hehe